OpenStudy (anonymous):

Suppose the population of a town is 500,000 in 1990. The population increases at a rate of 5% every five years. What will be the population of the town in 2005?

jimthompson5910 (jim_thompson5910):

We now use the formula A = P(1+r)^t where P is the initial population, r is the rate of increase, and t is the amount of time So now P = 500000, r = 0.05, and t = 15 (since 15 years elapse from 1990 to 2005)

jimthompson5910 (jim_thompson5910):

So we can say A = 500000(1+0.05)^15

OpenStudy (anonymous):

A= 500000(1.05)^15

jimthompson5910 (jim_thompson5910):

so far so good

jimthompson5910 (jim_thompson5910):

(1.05)^15 is what?

OpenStudy (anonymous):

2.07892818

jimthompson5910 (jim_thompson5910):

good, so that means A= 500000(1.05)^15 becomes A= 500000*2.07892818

jimthompson5910 (jim_thompson5910):

then multiply that to get ...?

OpenStudy (anonymous):

$1,039,464

jimthompson5910 (jim_thompson5910):

you have the right number (assuming you round to the nearest person), but no need for the $ sign (unless each person is a dollar lol)

OpenStudy (anonymous):

not to butt in, but it says 5% every five years. from 1990 to 2005 is 15 years, or 3 five year periods

jimthompson5910 (jim_thompson5910):

oh good point, didn't see that satellite, thx for pointing that out

OpenStudy (anonymous):

yw

jimthompson5910 (jim_thompson5910):

so it's really A = P(1+r)^(t/5)

jimthompson5910 (jim_thompson5910):

so this means that A = P(1+r)^(t/5) A = 500000(1+0.05)^(15/5) A = 500000(1.05)^(15/5) A = 500000(1.05)^3 A = 500000*1.157625 A = 578812.5 Telling us that the population will be about 578,812 people in the year 2005

OpenStudy (anonymous):

Wouln't it round to 578,813 because of the 0.5?

jimthompson5910 (jim_thompson5910):

well I rounded down because that 0.5 represents a person that's not really there

OpenStudy (anonymous):

Oh! Haha, thats right, we are talking about people =) Thank you!

jimthompson5910 (jim_thompson5910):

lol you're welcome